OpenStudy (destinyyyy):

Graph the following region-- 2x-5y>=0

OpenStudy (anonymous):

Ignoring the idea of regions, would you know how to graph 2x-5y = 0 as a line?

OpenStudy (destinyyyy):

No.. I could if zero was another number. I have no examples for this kind of problem

OpenStudy (anonymous):

The fact that the other number is 0 actually just means that you have a y-intersect of 0. Are you familiar with slope-intercept form?

OpenStudy (destinyyyy):

Yes

OpenStudy (anonymous):

Alright. So if I were to put your inequality into slope-intercept form I would have: \(2x - 5y \ge\ 0\) \(2x \ge\ 5y\) \(\frac{2}{5}x \ge\ y\) So that it maybe looks in a better order, let's rewrite this as \(y \le\ \frac{2}{5}x\) Now, you know slope-intercept form is \(y = mx+b\) where m is the slope and b is the y-intercept. So here we have a slope of 2/5, but no value for b. This means that the y-intercept is 0. So with that information, we can graph this line. Would you be able to graph it given the info above?

OpenStudy (anonymous):

Graph just the line I mean, don't worry about the region part yet.

OpenStudy (destinyyyy):

Yes?

OpenStudy (anonymous):

Just making sure :) Well, if we start from the y-intercept of 0 and use a slope of 2/5, we can get this line:  So how the line is drawn makes sense?